# Deterministic Dynamic Edge-Colouring

### v1

@@ 6.4 Analysing the data structures
Proof.
[…]
After colouring $e$, the only new vertices of $X_{1}$ that need to be added are representatives of bichromatic paths containing $e$. Any bichromatic component $C$ containing $e$ can only contain edges with colours $\kappa$ and $\kappa^{\prime}$ for some $\kappa^{\prime}\neq\kappa$. Since the colouring maintained is forced to be proper, it follows that $C$ must be either an even cycle or a path (possibly of length 0). In particular, $C$ can always be represented by a pseudo-tree or a tree with one additional special edge. Therefore, the algorithm always correctly identifies the number of of vertices in $C$ of degree $1$ or 00. Hence, the algorithm will always identify if $C$ is a (possibly empty) bichromatic path, and correctly insert a representative of $C$ into $\mathcal{B}(G)$. So $X_{1}$ contains a representative of all current bichromatic paths. The algorithm also deletes any bichromatic path whose Top-tree representation was modified during the process above. Since $e$ is inserted into the representation of $G[\kappa,\kappa^{\prime}]$ for all $\kappa^{\prime}$, either the Top-tree or the related special slot representing the component of $e$ in $G[\kappa,\kappa^{\prime}]$ will be modified. In particular, either the Top-tree or the related special slot representing bichromatic paths no longer present will be modified and hence deleted. Therefore, $X_{1}$ is maintained correctly.
@@
To recover the update times, we note that any operation to the Top-trees can be performed in $O(\log n)$ time (see [AHdLT05]), and that any non-local search for edges or degree $1$ or 00 vertices can be done in $O(\log n)$ time [AHdLT05] per edge. It is also standard that the number of such vertices can be maintained with only $O(1)$ overhead. Since we update or search only $O(\Delta_{\max})$ Top-trees, the total time spend searching in the Top-trees is in $O(\Delta_{\max}\cdot{}\ell\cdot{}\log n)$. Furthermore, for each new (and old) bichromatic path, of which there are at most $O(\Delta_{\max})$, we have to add (and delete) $2\Delta_{\max}$ edges to (from) $Y_{1}$ and $2\ell$ edges to (from) $Y_{3}$. Each edge is easy to insert (or delete) in constant time after performing the searches in the Top-trees already accounted for. Hence this takes additional time at most $O(\Delta_{\max}^{2}+\Delta_{max}\cdot{}\ell\cdot{}\log n)$ and the lemma follows. ∎
Remark 50.
@@
Lemma 52.
Assuming that all data-structures $\mathcal{D}_{i}$ are initially correct, then after the marking process of a call to $\emph{{colour}}(e,\kappa)$ or $\emph{{uncolour}}(e)$ concludes, we have that: if $f$ is any edge and $w$ any vertex with data-structure $\mathcal{D}_{i}(f,w)$ no longer correct, then $f$ is marked with $j\leq i$.
Proof.
@@
There are two reasons why $\mathcal{D}_{i}(f,w)$ might no longer be correct. The first reason $A)$ is that it contained an $i$-step process with step-length $\leq\ell$ which used a changed bichromatic path and so is no longer valid. Similarly it could also be that a new bichromatic path allows for a shorter consistent and augmenting $i$-step process with step-length $\leq\ell$, and this needs to be reflected in $\mathcal{D}_{i}(f,w)$. The second reason $B)$ is that changes to the colouring $c$ causes some invocations of Lemma 13 to be altered and so some of the stored processes are no longer consistent. Of course both reasons could be present at the same time. Observe that it is sufficient to show that if any $i$-step process with step-length $\leq\ell$ beginning with $w+x_{2}+x_{1}$ can reach either 1) a changed bichromatic path in $(i-1)$-steps or less or 2)2)2 ) any vertex in the one-hop neighbourhood of $e$ in $(i-1)$-steps or less, then $f$ is marked with $j\leq i$. Indeed, if 1)1)1 ) does not hold, then the reason $A)$ can never occur. Similarly, if 2)2)2 ) does not hold, then the reason $B)$ cannot occur.
@@
Thus it is sufficient to show that if 1)1)1 ) or 2)2)2 ) holds then $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ for some $2\leq j\leq i$ or $f$ is incident to a vertex in $R$. Indeed, then $f$ is marked with $j\leq i$.
@@
We begin by assuming that 1)1)1 ) holds. This means that after the changes to $c$, there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ using a changed bichromatic path $P$ during the $j^{\text{th}}$ step i.e. $P_{j}$ is a sub-path of $P$. Let $j\leq i$ be the smallest number for which the above is true.
@@
This means that before the change to $c$, $w_{1}+y_{1}+P_{1}+\dots+w_{j}+y_{j}$ was also an initial segment of some $j$-step process. Hence, this initial segment corresponds to a directed path in $\mathcal{B}(G)$ of the form $(r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{j}),r(y_{j})$ present both before and after the update to $c$. This follows from the minimality of $j$ and the discussion directly after Claim 48. Observe that then $y_{j}\in R$, $(r(w_{j})r(y_{j}))\in Y_{2}$, $(r(P_{j-1})r(w_{j}))\in Y_{3}$. Hence, if $j=1$, then $f=(x_{2}x_{1})=(r(y_{1})r(P_{1}))$ is incident to $R$, and so we would be done. More generally note that for all $k\leq j-1$, we have that $(r(y_{k})r(P_{k}))\in Y_{1}$, $(r(w_{k})r(y_{k}))\in Y_{2}$, and if $k>1$ that $(r(P_{k-1})r(w_{k}))\in Y_{3}$. Thus if $\psi:\mathcal{P}(V(\mathcal{B}(G)))\rightarrow\mathcal{P}(V(\mathcal{B}(G)))$ maps a set $A\in\mathcal{P}(V(\mathcal{B}(G)))$ to the set $N_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(A)))$, then $y_{k-1}\in\psi(\{y_{k}\})$ for $1<k\leq j$. In particular, it follows by induction that $y_{1}\in\psi^{j-1}(\{y_{j}\})$. So if $y_{j}\in R$, then $y_{1}\in M_{j}$ and therefore $f=(r(y_{1})r(P_{1}))\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$.
@@
On the other hand, assume that 2)2)2 ) holds. Then there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ such that for some $j$ we have $w_{j}=r(z)$ for some $z\in N^{1}(e)\cup e$ in the coloured inclusive neighbourhood of $z$. Indeed, otherwise the application of Lemma 13 cannot be changed at $z$, as it depends only on the colours of edges incident to vertices in the $1$-hop neighbourhood of $z$. Therefore there exist a vertex $z^{\prime}$ in the coloured neighbourhood of $z$ whose representative $r(z^{\prime})$ belongs to $R$. WLOG we can assume that $j$ is the smallest such index among all choices.
@@
If $j=1$, then we are now done as $r^{-1}(y_{1})$ must be one such vertex and therefore $f$ is incident to a vertex in $R$. For $j>1$ we can conclude, similarly to case 1)1)1 ), that $y_{1}\in\psi^{j-1}(R)$ and therfore that $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ and thusly marked with $j$. This concludes the proof of the lemma. ∎
Next we show that assuming that all un-marked data structures are maintained correctly, then we will update all marked structures correctly. Together with Lemma 52 above, this shows correctness of the update-algorithm.
Lemma 53.
[…]
Proof.
[…]
The base case is simple, as we just check all possible choices for semi-consistent and augmenting $1$-step processes with step-length at most $\ell$. At most one such process can exist, and if so, we store it. Note that this information can be computed without requiring information stored at any other data structures of the form $D_{\cdot}(\cdot,\cdot)$.
For the induction step, we assume that all data-structures $D_{k}(f,z)$ are correct for $k<i$. We fix choices of $f^{\prime}$ and $z^{\prime}$, and show that $D_{i}(f^{\prime},z^{\prime})$ is correctly computed. We first observe that the values $L(x)$ correctly lower bound the shortest semi-consistent and augmenting $i$-step process with step-length at most $\ell$ extended through $x$ at the $2^{\text{nd}}$ step.
Claim 54.
Let $f^{\prime}=(x_{2}x_{1}),z^{\prime}$ be as above. Let $x$ be any extension-point corresponding to some $P_{1}\in r^{-1}(x_{1})$. Then, at all times $L(x)$ is a lower-bound for the shortest semi-consistent and augmenting $i$-step process with step-length at most $\ell$ beginning with $z^{\prime}+r^{-1}(x_{2})+P_{1}+x$ not already present in $D_{i}(f^{\prime},z^{\prime})$.
Proof.
Any semi-consistent and augmenting $i$-step process with step-length at most $\ell$ beginning with $z^{\prime}+x_{2}+P_{1}+x$ must be extended by a semi-consistent and augmenting $(i-1)$-step process with step-length at most $\ell$ beginning with $x+y_{2}+P_{2}$ for some sub-path $P_{2}$ of $P$, where $y_{2}$ and $P$ are chosen consistently with Lemma 13, that furthermore does not overlap $z^{\prime}+x_{2}+P_{1}$. This is straight-forward to verify via Definition 15. By the induction hypothesis, we have that whenever $L(x)$ is set, it is set to $|P_{1}|+1+$ the length of the shortest semi-consistent and augmenting $(i-1)$-step process $C$ with step-length at most $\ell$ that satisfies the following:
  1. 1.
    $C$ begins with $x+y_{2}+P_{2}$ for some sub-path $P_{2}$ of $P$, where $y_{2}$ and $P$ are chosen consistently with Lemma 13.
  2. 2.
    $C$ is shortest among such processes that is not either already represented in $D_{i}(f^{\prime},z^{\prime})$ or has already been confirmed to overlap $z^{\prime}+x_{2}+P_{1}$ in the sense of Definition 15.
In particular, it is clear that $L(x)$ therefore lower bounds the length of any sought-after process. ∎
Secondly, we observe that if we have $\alpha$ semi-consistent and augmenting $(i-1)$-step processes with step-length at most $\ell$ beginning with $x+y_{2}+P_{2}$ for some $P_{2}$ with the choices of $y_{2}$ and $P_{2}$ consistent with Lemma 13, then $D_{i-1}((r(y_{2})r(P_{2}),r(x))$ contains at least $\min\{\alpha,\sqrt{\log n}(\sqrt{\log n}-(i-1))\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell\}$ augmenting and semi-consistent $(i-1)$-step processes which can be added to the $1$-step process $z+r^{-1}(x_{2})+P_{1}$ to form either all or the shortest $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ semi-consistent and augmenting processes of the form we seek.
Indeed, we note that by applying Lemma 41 to $z^{\prime}+r^{-1}(x_{2})+P_{1}$, we find that at most $i\cdot{}3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}\ell$ of all non-overlapping $j$-step processes with $j\leq i-1$ can overlap with $z^{\prime}+r^{-1}(x_{2})+P_{1}$ in the sense of Definition 15. Note that this bound also holds for the extensions stored in $D_{i-1}((r(y_{2})r(P_{2}),r(x))$, as the first path of any extension can only overlap the $2$-hop neighbourhood of $z^{\prime}$, in which case the extension is accounted for above via Lemma 41, or via a later part of the process that loops back and overlaps the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$. But this later part of the process then forms a $j$-step process for $j<i-1$ that starts outside of the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$ and then reaches the two hop-neighbourhood of $z^{\prime}+r^{-1}(x_{2})+P_{1}$. Thus it is also accounted for by Lemma 41.
In particular, since $D_{i-1}((r(y_{2})r(P_{2}),r(x))$ contains the shortest
$\sqrt{\log n}(\sqrt{\log n}-(i-1))\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$
candidate extensions (or fewer, if fewer exists), we either find all extensions that will not overlap with $z^{\prime}+r^{-1}(x_{2})+P_{1}$ in the sense of Definition 15 or we find $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ such extensions. In particular, the above together with Claim 54 implies that the algorithm will locate either all or the $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$ shortest semi-consistent and augmenting $i$-step processes with step-length at most $\ell$ beginning with $z^{\prime}+r^{-1}(x_{2})+P^{\prime}$, where $P^{\prime}\in r^{-1}(x_{1})$. This concludes the induction step. ∎
Lemma 53, Lemma 52, and Lemma 49 together show correctness of the data structure and the update algorithms for the supported operations.
@@
We first briefly comment on implementations certain data structures. For each BBST, we can use Red-blacks trees [CLRS22] to support basic operations in $O(\log n)$ worst-case time. For Top-trees we use the version from [AHdLT05]. As for the analysis, we first show that not too many edges are marked:
Lemma 55.
@@
Proof.
The proof is by induction on $i$. For $i=1$, we note that by item two and item three of Claim 48 at most $4(\Delta_{\max}+1)^{2}\cdot{}|R|$ edges are initially marked $1$. Also $|M_{1}|=|R|$. For the induction step, we assume the statement for all $j\leq i-1$. We note that applying item two of Claim 48 thrice shows that $M_{i}\leq 2^{3}(\Delta_{\max}+1)^{6}|M_{i-1}|$ and that the number of edges marked with $i$ is also at most $2^{2}(\Delta_{\max}+1)^{6}|M_{i-1}|$. Applying the induction hypothesis now yields the result. ∎
@@
Lemma 56.
In each invocation of the marking scheme, we have that $|R|\leq 2+2(\Delta_{\max}+1)+6(\Delta_{\max}^{2}+1)^{2}\leq 10(\Delta_{\max}^{2}+1)^{2}$.
Proof.
We note that at most $3(\Delta_{\max}^{2}+1)^{2}$ bichromatic paths are altered after a removal or insertion of a coloured edge: namely 3 for each pair of colours; the bichromatic path that used to contain the edge and the two new bichromatic paths that this path was split into (for insertion it is of course opposite; one new bichromatic path that is the fusion of two old). There are at most $(\Delta_{\max}+1)^{2}$ choices for such colour pairs. Note that each path has at most $2$ endpoints. Finally, there are at most $2(\Delta_{\max}+1)$ vertices in the one-hop neighbourhood of $u$ and $v$. Adding these quantities gives the bound. ∎
@@
Proof.
We can mark each edge with constant overhead by performing a backwards BFS search along the relevant edges (assuming we store different types of edges in different BBST’s). Hence, the Lemma follows by combining Lemma 55 with Lemma 56, and summing over all choices of $1\leq i\leq\sqrt{\log n}$:
@@
Lemma 58.
Given a $1$-step process $w_{1}+y_{1}+P_{1}$, with $r^{-1}(w_{2})$ the second last vertex of $P_{1}$, and a semi-consistent and augmenting $(i-1)$-step process $w_{2}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$, we can check if $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ is a semi-consistent and augmenting $i$-step process in $O((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time.
Proof.
We check whether $w_{2},y_{2}$ and $P_{2}$ is consistent with Lemma 13 in $\tilde{O}(\Delta_{\max}^{3})$. Next we can can trace and temporarily mark the two-hop neighbourhood of every vertex in $w_{1}+y_{1}+P_{1}$, as well as the edges in $P_{1}$. Finally, we can follow along the $(i-1)$-step process and check the conditions of Definition 15 by using the temporary marks. It takes $\tilde{O}((i+1)\ell)$ time to trace the process and $\tilde{O}(\ell(\Delta_{\max}+1)^{2})$ to temporarily mark all of the edges and vertices. We note here that any update time required to correctly locate edges or vertices in memory is captured by the $\tilde{O}$-notation which accounts for look-ups in BBST. ∎
We have already seen in the proof of Lemma 53 that not too many of the processes stored in neighbouring data structures can be overlapping. This is captured in the following Lemma:
Lemma 59.
Suppose we are given a $1$-step process $w_{1}+y_{1}+P_{1}$, with $r^{-1}(w_{2})$ the second last vertex of $P_{1}$. Let $C^{\prime}$ be a semi-consistent and augmenting $(i-1)$-step processes of the form $C^{\prime}=w_{2}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$. Suppose further that the choice of $w_{2},y_{2}$ and $P_{2}$ is consistent with Lemma 13. Then at most $i\cdot{}3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}\ell$ choices of $C^{\prime}$ can cause $C=w_{1}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$ to not be a semi-consistent and augmenting $i$-step process.
Proof.
Since the extension is already assumed to be consistent with Lemma 13, we need only check that $C^{\prime}$ does not overlap $w_{1}+y_{1}+P_{1}$ in the sense of Definition 15. We can however bound the number of such $C^{\prime}$ reaching the two-hop neighbourhood of $w_{1}$ by applying Corollary 42 to every point in this neighbourhood. We note that due to the extension being consistent with Lemma 13, any other violations has to be because $P_{j}$ intersects the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$ for some $j\geq 3$ or because we reuse the maximal bichromatic path containing $P_{1}$. In the first-case the non-overlapping process $C^{\prime}=w_{3}+y_{3}+P_{3}+\dots+w_{i}+y_{i}+P_{i}$ also intersects the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$. But since $w_{3}$ is disjoint from the $2$-hop neighbourhood of $w_{1}+y_{1}+P_{1}$, this process is accounted for by Corollary 42. The second case is similarly accounted for by Corollary 42. ∎

### v2

@@ 6.4 Analysing the data structures
Proof 31.
We have already described the algorithm in Section 6.3, so what remains is to analyse it and prove that it is correct. To show correctness, we have to prove two things. First of all that the maintained bichromatic skeleton has the required vertex set, and second of all that it has the required edge-set. Since $X_{2}$ cannot be updated, we need only check that $X_{1}$ is correctly maintained. Since $\texttt{add}(e)$ and $\texttt{remove}(e)$ does not alter $X_{1}$, we need only check that $\texttt{un-colour}(e)$ and $\texttt{colour}(e,\kappa)$ correctly re-establishes $X_{1}$. We will argue that $\texttt{colour}(e,\kappa)$ does so, and note that the case for $\texttt{un-colour}(e)$ is completely symmetrical.
After colouring $e$, the only new vertices of $X_{1}$ that need to be added are representatives of bichromatic paths containing $e$. Any bichromatic component $C$ containing $e$ can only contain edges with colours $\kappa$ and $\kappa^{\prime}$ for some $\kappa^{\prime}\neq\kappa$. Since the colouring maintained is forced to be proper, it follows that $C$ must be either an even cycle or a path (possibly of length 0). In particular, $C$ can always be represented by a pseudo-tree or a tree with one additional special edge. Therefore, the algorithm always correctly identifies the number of of vertices in $C$ of degree $1$ or 0. Hence, the algorithm will always identify if $C$ is a (possibly empty) bichromatic path, and correctly insert a representative of $C$ into $\mathcal{B}(G)$. So $X_{1}$ contains a representative of all current bichromatic paths. The algorithm also deletes any bichromatic path whose Top-tree representation was modified during the process above. Since $e$ is inserted into the representation of $G[\kappa,\kappa^{\prime}]$ for all $\kappa^{\prime}$, either the Top-tree or the related special slot representing the component of $e$ in $G[\kappa,\kappa^{\prime}]$ will be modified. In particular, either the Top-tree or the related special slot representing bichromatic paths no longer present will be modified and hence deleted. Therefore, $X_{1}$ is maintained correctly.
@@
To recover the update times, we note that any operation to the Top-trees can be performed in $O(\log n)$ time (see [HolmEtAL]), and that any non-local search for edges or degree $1$ or 0 vertices can be done in $O(\log n)$ time [HolmEtAL] per edge. It is also standard that the number of such vertices can be maintained with only $O(1)$ overhead. Since we update or search only $O(\Delta_{\max})$ Top-trees, the total time spend searching in the Top-trees is in $O(\Delta_{\max}\cdot{}\ell\cdot{}\log n)$. Furthermore, for each new (and old) bichromatic path, of which there are at most $O(\Delta_{\max})$, we have to add (and delete) $2(\Delta_{\max}+1)$ edges to (from) $Y_{1}$ and $2\ell$ edges to (from) $Y_{3}$. Each edge is easy to insert (or delete) in constant time after performing the searches in the Top-trees already accounted for. Hence this takes additional time at most $O(\Delta_{\max}^{2}+\Delta_{\max}\cdot{}\ell\cdot{}\log n)$ and the lemma follows.
Remark 6.22.
@@
The algorithm needs two parameters $\ell$ and $a$, which we will specify later. The only important thing is that they satisfy the requirements of earlier presented lemmas. We will assume this is the case for now, and then fix the parameters later.
@@
Lemma 6.24.
Assuming that all data structures $\mathcal{D}_{i}$ are initially valid, then after the marking process of a call to $\emph{{colour}}(e,\kappa)$ or $\emph{{uncolour}}(e)$ concludes, we have that: if $f$ is any edge and $w$ any vertex with data structure $\mathcal{D}_{i}(f,w)$ no longer valid, then $f$ is marked with $j\leq i$.
Proof 32.
@@
There are two reasons why $\mathcal{D}_{i}(f,w)$ might no longer be correct. The first reason $A)$ is that it contained an $i$-step process with step-length $\leq\ell$ which used a changed bichromatic path and so is no longer valid. The second reason $B)$ is that changes to the colouring $c$ causes some invocations of Lemma 3.5 to be altered and so some of the stored processes are no longer semi-consistent. Of course both reasons could be present at the same time. Observe that it is sufficient to show that if any $i$-step process with step-length $\leq\ell$ beginning with $w+x_{2}+x_{1}$ can reach either 1) a changed bichromatic path in (i1)(i-1)-steps or less or $2)$ any vertex in the one-hop neighbourhood of $e$ in (i1)(i-1)-steps or less, then $f$ is marked with $j\leq i$. Indeed, if $1)$ does not hold, then the reason $A)$ can never occur. Similarly, if $2)$ does not hold, then reason $B)$ cannot occur.
@@
Thus it is sufficient to show that if $1)$ or $2)$ holds then $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ for some $2\leq j\leq i$ or $f$ is incident to a vertex in $R$. Indeed, then $f$ is marked with $j\leq i$.
@@
We begin by assuming that $1)$ holds. This means that after the changes to $c$, there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ using a changed bichromatic path $P$ during the $j^{\text{th}}$ step i.e. $P_{j}$ is a sub-path of $P$. Let $j\leq i$ be the smallest number for which the above is true.
@@
This means that before the change to $c$, $w_{1}+y_{1}+P_{1}+\dots+w_{j}+y_{j}$ was also an initial segment of some $j$-step process. Hence, this initial segment corresponds to a directed path in $\mathcal{B}(G)$ of the form $r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{j}),r(y_{j})$ present both before and after the update to $c$. This follows from the minimality of $j$ and the discussion directly after Claim 6.20. Observe that then $y_{j}\in R$, $(r(w_{j})r(y_{j}))\in Y_{2}$, $(r(P_{j-1})r(w_{j}))\in Y_{3}$. Hence, if $j=1$, then $f=(x_{2}x_{1})=(r(y_{1})r(P_{1}))$ is incident to $R$, and so we would be done. More generally note that for all $k\leq j-1$, we have that $(r(y_{k})r(P_{k}))\in Y_{1}$, $(r(w_{k})r(y_{k}))\in Y_{2}$, and if $k>1$ that $(r(P_{k-1})r(w_{k}))\in Y_{3}$. Thus if $\psi:\mathcal{P}(V(\mathcal{B}(G)))\rightarrow\mathcal{P}(V(\mathcal{B}(G)))$ maps a set $A\in\mathcal{P}(V(\mathcal{B}(G)))$ to the set $N_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(A)))$, then $r(y_{k-1})\in\psi(\{r(y_{k})\})$ for $1<k\leq j$. In particular, it follows by induction that $y_{1}\in\psi^{j-1}(\{y_{j}\})$. So if $y_{j}\in R$, then $y_{1}\in M_{j}$ and therefore $f=(r(y_{1})r(P_{1}))\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$.
@@
On the other hand, assume that $2)$ holds. Then there is an $i$-step process $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ with $w_{1}=w$, $y_{1}=r^{-1}(x_{2})$, and $P_{1}\in r^{-1}(x_{1})$ such that for some $j$ we have $w_{j}=z$ for some $z\in N^{1}(e)$ in the coloured neighbourhood of $e$. I.e. we only consider neighbours of $z$ along edges that are coloured. Indeed, otherwise the application of Lemma 3.5 cannot be changed at $z$, as it depends only on the colours of edges incident to vertices in the $1$-hop neighbourhood of $z$. WLOG we can assume that $j$ is the smallest such index among all choices.
@@
If $j=1$, then we are now done as $y_{1}\in N^{1}(z)\subset N^{2}(e)$, and therefore $f$ is incident to a vertex in $R$. For $j>1$ we can conclude, similarly to case $1)$, that $r(y_{1})\in\psi^{j-1}(R)$ and therfore that $f\in E_{Y_{1},\text{in}}(N_{Y_{3},\text{in}}(N_{Y_{2},\text{in}}(M_{j-1})))$ and thusly marked with $j$. This concludes the proof of the lemma.
The algorithm maintain sets $S^{i}_{w,y,P}$ for all $i\leq\sqrt{\log n}$ in accordance with the description before Corollary 6.13, so since the lemma above together with induction shows that these sets are properly updated, we can apply Corollary 6.13. Hence, we conclude that for any valid choice of $w,y$, and $P$, we find that the set $S^{\sqrt{\log n}}_{w,y,P}$ will be non-empty. Finally, Lemma 6.12 and the subsequent discussion proves that if $w$ is $(a,i,\emptyset,\emptyset,P)$-heavy, then the set $S^{i}_{w,y,P}$ is non-empty. This concludes correctness of the algorithm.
In the following paragraph, we provide implementations of the data structures used by the algorithm and analyse the update time.
@@
We first briefly comment on implementations certain data structures. For each BBST, we can use Red-blacks trees [cormen2022introduction] to support basic operations in $O(\log n)$ worst-case time. For Top-trees we use the version from [HolmEtAL]. As for the analysis, we first show that not too many edges are marked:
Lemma 6.25.
@@
Proof 33.
The proof is by induction on $i$. For $i=1$, we note that by item two and item three of Claim 6.20 at most $4(\Delta_{\max}+1)^{2}\cdot{}|R|$ edges are initially marked $1$. Also $|M_{1}|=|R|$. For the induction step, we assume the statement for all $j\leq i-1$. We note that applying item two of Claim 6.20 thrice shows that $M_{i}\leq 2^{3}(\Delta_{\max}+1)^{6}|M_{i-1}|$ and that the number of edges marked with $i$ is also at most $2^{3}(\Delta_{\max}+1)^{6}|M_{i-1}|$. Applying the induction hypothesis now yields the result.
@@
Lemma 6.26.
In each invocation of the marking scheme, we have that $|R|\leq 2+2(\Delta_{\max}+1)^{2}+6(\Delta_{\max}^{2}+1)^{2}\leq 10(\Delta_{\max}^{2}+1)^{2}$.
Proof 34.
We note that at most $3(\Delta_{\max}^{2}+1)^{2}$ bichromatic paths are altered after a removal or insertion of a coloured edge: namely 3 for each pair of colours; the bichromatic path that used to contain the edge and the two new bichromatic paths that this path was split into (for insertion it is of course opposite; one new bichromatic path that is the fusion of two old). There are at most $(\Delta_{\max}+1)^{2}$ choices for such colour pairs. Note that each path has at most $2$ endpoints. Finally, there are at most $2(\Delta_{\max}+1)^{2}$ vertices in the $2$-hop neighbourhood of $u$ and $v$. Adding these quantities gives the bound.
@@
Proof 35.
We can mark each edge with constant overhead by performing a backwards BFS search along the relevant edges (assuming we store different types of edges in different BBST’s). Hence, the Lemma follows by combining Lemma 6.25 with Lemma 6.26, and summing over all choices of $1\leq i\leq\sqrt{\log n}$:
@@
Lemma 6.28.
Given a $1$-step process $w_{1}+y_{1}+P_{1}$, with $w_{2}$ the last vertex of $P_{1}$, and a semi-consistent and augmenting (i1)(i-1)-step process $w_{2}+y_{2}+P_{2}+\dots+w_{i}+y_{i}+P_{i}$, we can check if $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ is a semi-consistent and augmenting $i$-step process in $\tilde{O}((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time.
Proof 36.
We check whether $w_{2},y_{2}$ and $P_{2}$ is consistent with Lemma 3.5 in $\tilde{O}(\Delta_{\max}^{3})$ time. Next we can can trace and temporarily mark the two-hop neighbourhood of every vertex in $w_{1}+y_{1}+P_{1}$, as well as the edges in $P_{1}$. Finally, we can follow along the (i1)(i-1)-step process and check the conditions of Definition 3.8 by using the temporary marks. It takes $\tilde{O}((i+1)\ell)$ time to trace the process and $\tilde{O}(\ell(\Delta_{\max}+1)^{2})$ to temporarily mark all of the edges and vertices. We note here that any update time required to correctly locate edges or vertices in memory is captured by the $\tilde{O}-$notation which accounts for look-ups in the BBST. We can use the Top-trees to check whether a subpath from a previously used maximal bichromatic path is present.
In particular, the lemma above essentially says that we can check if a process is blocked in $\tilde{O}((i+1)\cdot{}\ell\cdot{}(\Delta_{\max}+1)^{3})$ time. Indeed, checks not accounted for above can be done in a similar fashion. Hence, we have the following lemma:
Lemma 6.29.
The algorithm can check if a particular extension point $x$ is blocked in $\tilde{O}(\sqrt{\log n}\cdot{}\ell\cdot{}a^{2}\cdot{}(\Delta_{\max}+1)^{3})$ time. In addition, it can mark any non-blocked strings with $O(1)$ overhead.
Proof 37.
For any choice of $x$, we need only check if it is blocked at 3 different levels. In total at most $O(a^{2})$ process represented by a string from the relevant set needs to be checked. Each process has at most $\sqrt{\log n}$ steps, and the lemma follows by applying Lemma 6.28.